# Topological Spatial Graph Coarsening

### v1

@@ Abstract
Spatial graphs are particular graphs for which the nodes are localized in space (e.g., public transport network, molecules, branching biological structures). In this work, we consider the problem of spatial graph reduction, that aims to find a smaller spatial graph (i.e., with less nodes) with the same overall structure as the initial one. In this context, performing the graph reduction while preserving the main topological features of the initial graph is particularly relevant, due to the additional spatial information. Thus, we propose a topological spatial graph coarsening approach based on a new framework that finds a trade-off between the graph reduction and the preservation of the topological characteristics. The coarsening is realized by collapsing short edges. In order to capture the topological information required to calibrate the reduction level, we adapt the construction of classical topological descriptors made for point clouds (the so-called persistent diagrams) to spatial graphs. This construction relies on the introduction of a new filtration called triangle-aware graph filtration. Our coarsening approach is parameter-free and we prove that it is equivariant under rotations, translations and scaling of the initial spatial graph. We evaluate the performances of our method on synthetic and real spatial graphs, and show that it significantly reduces the graph sizes while preserving the relevant topological information.
@@
Spatial graphs, Graph reduction, Coarsening, Topological data analysis
@@ 1 Introduction
Spatial graphs are a specific type of graphs with spatial attributes associated to the nodes and edges. They are studied in different applicative context and called with various names. In transport analysis, social network analysis, and epidemiology, they are often referred to as spatial networks [barthelemy2011spatial, gou2021understanding, doytsher2010querying]. Spatial graphs are also closely linked to geometric graphs — which are graphs whose nodes lie in Euclidean space and whose edges are (possibly intersecting) straight lines [cerny2005geometric]. When edges may instead be continuous curves, the structure is known as a topological or metric graph [pach1997graphs]. Geometric, topological, and metric graphs are extensively studied within graph theory, which explores the possible configurations of a graph with a given number of nodes and constraints on edges; see [pach2013beginnings, pach2004geometric] for historical and technical perspectives. Finally, spatial graphs can also be viewed as simplified forms of shape or elastic graphs, where actual shapes are encoded along the edges [guo2020representations].
When working with such graph, a common problem is dimensionality as the graphs have a high number of nodes and edges. Airways in the lungs or neurons networks are common real world examples displaying high dimensionality. The problem of decreasing the graph complexity has been studied in the network analysis literature under the name of graph sparsification or graph coarsening (see [hashemi2024comprehensive] for a recent overview of the field). Graph sparsification consists in selecting only a subset of nodes and edges [althofer1993sparse], while graph coarsening consists in grouping the nodes into super-nodes and edges into super-edges using aggregation algorithms [loukas2018spectrally].
In this paper, we propose a way to simplify a spatial graph by preserving its topology. We want to preserve the principal graph structure (connected components, large cycles) while removing small and noisy graph components. The principal adjective recalls the principal component analysis as our method is similarly trying to capture the characteristics of the graph which retain the majority of the information. We propose a spatial graph coarsening procedure which is topologically informed - here called Topological Spatial Graph Coarsening. To the best of our knowledge, the proposed method is new in the literature and addresses the open challenge in many real world applications of working with graphs with lower number of nodes.
@@
To define a dimensionality reduction procedure which is topologically informed, we need to estimate the topological properties of the spatial graphs. We leverage tools from Topological Data Analysis (TDA) [chazal2021introduction]. Taking its root in the field of algebraic topology, TDA is a recent and promising domain that provides well-founded tools to analyze complex data - such as graphs - for which topological properties represent key insights into understanding the data. Here, by topology we refer to the properties left unchanged by continuous deformations such as rotation, translation, stretching, and bending (e.g., number of connected components or holes). A common tool in TDA is the persistent diagram, which is classically used to study the topology of point clouds in $\mathbb{R}^{p}$ - see seminal works by [edelsbrunner2002topological, carlsson2009topology, aktas2019persistence]. As for point clouds, some attempt have been done to adapt classic filtration to graph data [archambault2007topolayout, li2012effective, ferrara2011topological]. For example, a graph filtration can be defined based on the node degree [hofer2017deep], the Jaccard index and Ricci curvature [zhao2019learning], or the Heat Kernel [carriere2020perslay, lasalle2024heat]. As our graphs have a spatial component, we define a new filtration which is based on both the spatial and the structural information. We name it the triangle-aware graph filtration as it captures cycles information generated by all cycles, even the one induced by triangles in the graph.
Given the topological characterization, we can now propose a coarsening procedure for spatial graphs. The new graph preserve the most important topological features of the original graph -measured via our filtration and persistent diagram - with a lower number of nodes and edges. As for any dimensionality reduction technique, we would like to calibrate the reduction level, as too much reduction would yield to an oversimplified graph with missing important features, and too little reduction would not provide a significant and useful coarsening. We define a score to guide the choice of an appropriate reduction level. This score balances the reduced graph complexity with the amount of topological “distortion” between the original and reduced graph evaluated by comparing persistence diagrams. Minimizing this score allows to calibrate the reduction level.
@@
The paper is organized as follow. In Section 2, we introduce the spatial graph object and the coarsening procedure. In Section 3, we recall how to compute classical persistence diagrams and we introduce the persistence diagram adaptation for spatial graphs (Subsection 3.2). In Section 4, we define the topologically informed spatial graph coarsening. As the new methodology is tailored for spatial graphs, we study the invariance and equivariance properties of the method with respect to rotation, translation and scaling of the node positions; see Section 5. In Section 6, we showcase the proposed method on simulated data and on a real world dataset of fungi networks [fricker2025fungi].
@@ 2 Spatial Graph Coarsening
Consider a spatial graph $G$ consisting of a set of nodes $V=\{v_{1},\dots,v_{n}\}$ and a set of edges $E\in V\times V$. In a spatial graph, each node is equipped with spatial positions $(X_{u})_{u\in V}$, where $x_{u}\in\mathbb{R}^{p}$ for all nodes $u$, with $p$ being the dimension of the space. The spatial graph is denoted by the triplet $G=(V,E,X)$. In this work, the nodes are embedded in $\mathbb{R}^{p}$ where $p=2,3$ and equipped with Euclidean distance. For each edge $(u,v)\in E$, the length function $\ell:E\rightarrow\mathbb{R}_{+}$ assigns the edge length $\ell_{u,v}=\|x_{u}-x_{v}\|_{p}$.
The objective of this paper is to define a coarsening procedure for spatial graphs, retaining the main topological properties of the original graph. In a general graph setting, given a set of graphs $\mathcal{G}$, a coarsening method can be defined via a coarsening function $f:\mathcal{G}\rightarrow\mathcal{G}$ which defines a new graph on a partition of the nodes of the original graph: $f(G)=(f^{V}(G),f^{E}(G))$ where $f^{V}(G)\in\mathcal{P}(V)$, with $\mathcal{P}(V)$ denoted the set of partitions of $V$. As our graphs are spatial, we define a spatial graph coarsening methods based on a spatial parameter called $\theta$. Such parametrization of the coarsening procedure allows to define a coarsened graph which depends on both the structural and the spatial structure of the graph. Essentially, we obtain the coarsened spatial graph by collapsing the edges with a length smaller than $\theta$. Before giving the formal definition of our spatial graph coarsening in Definition 1, we introduce the notation $sub(G,\theta)$ that denotes the abstract subgraph of $G$ induced by the edges of length smaller than $\theta$. So, $sub(G,\theta)=(V,E_{\theta})$, such that $E_{\theta}=\{(u,v)\in E,\ s.t.\ \ell_{u,v}\leq\theta\}$.
@@
Consider a spatial graph $G=(V,E,X)$ and a parameter $\theta\in\mathbb{R}_{+}$. We define the coarsened graph at scale $\theta$ by $f_{\theta}(G)=(f_{\theta}^{V}(G),f_{\theta}^{E}(G),f_{\theta}^{X}(G))$ where the three components $f_{\theta}^{V}(G)$, $f_{\theta}^{E}(G)$ and $f_{\theta}^{X}(G))$, acting respectively on the nodes, edges and positions, are defined as follows:
@@
    • the average positioning assigns a new spatial position to a node $V_{i}\in f_{\theta}^{V}(G)$ as the average position of the merged nodes:
@@
    • the degree positioning is based on selecting the spatial position of the node with higher degree among the merged nodes:
@@
      where $deg(v)\in\mathbb{R}_{+}$ is the sum the weights of the incident edges in $v$.
@@
The parameter $\theta$ can take values in $\mathbb{R}_{+}$. However, the interesting set of values for the parameter when working in practice are $\Theta=[\min_{(u,v)\in E}\ell_{u,v},\max_{(u,v)\in E}\ell_{u,v}]$. The coarsed graph for $\theta>\max_{(u,v)\in E}\ell_{u,v}$ is the trivial graph where all the nodes are collapsed into one single node. To simplify the notation, we refer to $\theta\in\mathbb{R}_{+}$
Once we have introduced a spatial coarsening procedure, the question is now how to select the parameter $\theta$. The objective of the paper is to define a coarsened graphs with lower number of nodes while preserving the key topological features of the original graph. An illustrative example is shown in Figure 1. We sample $20$ nodes at random with position on an annulus shape. The corresponding fully connected graph has $(n^{2}-n)/2$ edges, but we select only a percentage of the edges, keeping only those with smallest length. The original graph is reduced into three graphs with lower number of nodes, obtained with different parameter $\theta$. From the picture it is clear to see how the coarsening procedure is first merging the nodes while preserving its key topological feature which is largest the cycle. As the parameter $\theta$ increases, more nodes are merged together, resulting in an oversimplified graph. To identify a parameter $\theta$ that performs a good balance between reducing the graph and preserving the main topological features, we define a procedure that relies on persistence diagrams to capture the topological information.
@@ 3 Persistence Diagrams for Spatial Graphs
Classical TDA pipelines consist in capturing certain properties for shapes underlying the data such as number of holes or number of connected components. One of the main tools is a topological descriptor called persistence diagram. In this section, we briefly introduce the general TDA pipeline for analyzing point clouds, before describing how they can be adapted to graph, and especially spatial graphs. For a more complete presentation of TDA, we refer the reader to [edelsbrunner2010computational, oudot2015persistence, chazal2021introduction, iniesta2022topological].
3.1 Classical persistence diagrams
Data as point clouds have no intrinsic interesting topology (only isolated contractible connected components). However, there often exist underlying continuous structures with richer topology, e.g., points sampled on some manifolds.
@@ 3.1 Classical persistence diagrams
To capture the topological invariants of the underlying shape, we construct continuous spaces on top of the point clouds called simplicial complexes. A simplicial complex is a space made by gluing together simple structures of different dimensions: points, segments, triangles, tetrahedrons etc. The simple building blocs are called simplices. They are represented in an abstract way as a set of their vertices, e.g., an segment between point $u$ and point $v$ is encoded as $\{u,v\}$, a triangle between $u$, $v$ and $w$ is given by $\{u,v,w\}$. Simplicial complexes are a generalization of classical graphs, which are one-dimensional objects (composed of points and edges only), where one can include higher-dimensional objects like triangles (dimension 2), tetrahedron (dimension 3) and equivalents in higher dimensions. Formally, they are a set of simplices, i.e., a set of sets of vertices. Given a point clouds $\mathcal{X}=\{x_{1},\dots x_{n}\}$ in $\mathbb{R}^{p}$ equipped with the Euclidean distance $d$, a classical simplicial complex built on top of $\mathcal{X}$ is the so-called Vietoris-Rips (VR) complex at scale $r>0$ denoted by $VR_{r}(\mathcal{X},d)$ [zomorodian2010fast]. To determine which simplices are included in $VR_{r}(\mathcal{X})$ consider all balls of radius $r$ centered on the points of $\mathcal{X}$. If two balls intersect, the corresponding edge between the two centers is an element of $VR_{r}(\mathcal{X},d)$. More generally, for higher dimension, any $k$-dimensional simplex (subset of $k+1$ points of $\mathcal{X}$) is included if all pairs of balls centered in points of the simplex intersect. Formally, the $k$-simplex $\{x_{i_{1}},\dots,x_{i_{k}}\}$ is in $VR_{r}(\mathcal{X},d)$, if $\forall\quad 1\leq j<j^{\prime}\leq k$, $d(x_{i_{j}},x_{i_{j^{\prime}}})\leq r$.
@@
From this continuous structures one can now compute topological invariants in the data [chazal2021introduction]. To extract meaningful topological information from the data, instead of looking at the topological invariants of one simplicial complex at a given scale $r>0$, we can look at their evolution in the family of simplicial complexes $(VR_{r}(\mathcal{X},d))_{r>0}$ as $r$ increases. Such a family is called a filtration, i.e., a family of nested simplicial complexes. As $r$ increases, some topological features (connected components, holes, cavities, etc.) appear and disappear. For each topological feature, we record the scale $r_{b}$ at which it appears (it is born) and the scale $r_{d}\geq r_{b}$ at which it disappears (it dies). We encode this information into a point $(r_{b},r_{d})$ in $\mathbb{R}^{2}$. The collection of such points for all topological features is called the persistence diagram (PD). Points that are closer to the diagonal $y=x$ in the PD corresponds to topological features that lived during a short range of scales, they did not persist, hence they are usually associated to topological noise. Conversely, points that lie far from the diagonal correspond to topological features that persisted in the filtration and therefore correspond to important topological characteristics of the data. Note that monitoring the topological invariants in the filtration avoids having to choose a specific scale at which relevant topological features would be present. How much a topological feature persists in the filtration encode its relevance as a topological information of the data.

### v2

@@ Abstract
Spatial graphs are particular graphs for which the nodes are localized in space (e.g., public transport network, molecules, branching biological structures). In this work, we consider the problem of spatial graph reduction, that aims to find a smaller spatial graph (i.e., with less nodes) with the same overall structure as the initial one. In this context, performing the graph reduction while preserving the main topological features of the initial graph is particularly relevant, due to the additional spatial information. Thus, we propose a topological spatial graph coarsening approach based on a new framework that finds a trade-off between the graph reduction and the preservation of the topological characteristics. In order to capture the topological information required to calibrate the reduction level, we adapt the construction of classic topological descriptors made for point clouds (the so-called persistent diagrams) to spatial graphs. This construction relies on the introduction of a new filtration called triangle-aware graph filtration. Our coarsening approach is parameter-free and we prove that it is equivariant under rotations, translations and scaling of the initial spatial graph. We test the method on synthetic and real spatial graphs (road network and fungi network), and show that it significantly reduces the graph sizes while preserving the relevant topological information.
@@
Spatial graphs, Graph Coarsening, Topological Data Analysis
@@ 1 Introduction
Spatial graphs are a specific type of graphs with spatial attributes associated to the nodes and edges. They are studied in different applicative context and called with various names. In transport analysis, social network analysis, and epidemiology, they are often referred to as spatial networks [barthelemy2011spatial, gou2021understanding, doytsher2010querying]. Spatial graphs are also closely linked to geometric graphs — which are graphs whose nodes lie in Euclidean space and whose edges are (possibly intersecting) straight lines [cerny2005geometric]. When edges may instead be continuous curves, the structure is known as a topological or metric graph [pach1997graphs]. Geometric, topological, and metric graphs are extensively studied within graph theory, which explores the possible configurations of a graph with a given number of nodes and constraints on edges; see [pach2013beginnings, pach2004geometric] for historical and technical perspectives. Finally, spatial graphs can also be viewed as simplified forms of shape or elastic graphs, where actual shapes are encoded along the edges [guo2020representations]. When working with such graphs, a common problem is dimensionality as they may have a high number of nodes and edges. Airways in the lungs or neurons networks are common real world examples displaying high dimensionality. The problem of decreasing the graph complexity has been studied in the network analysis literature under the name of graph sparsification or graph coarsenin, see [hashemi2024comprehensive] for a recent overview of the field. Graph sparsification consists in selecting only a subset of nodes and edges [althofer1993sparse], while graph coarsening consists in grouping the nodes into super-nodes and edges into super-edges using aggregation algorithms [loukas2018spectrally].
@@
In this paper, we propose a way to simplify a spatial graph by preserving its topology. We want to preserve the principal graph structure (connected components, large cycles) while removing small and noisy graph components. The principal adjective recalls the principal component analysis as our method is similarly trying to capture the characteristics of the graph which retain the majority of the information. We propose a spatial graph coarsening procedure which is topologically informed - here called Topological Spatial Graph Coarsening. To the best of our knowledge, the proposed method is new in the literature and addresses the open challenge in many real world applications of working with graphs with lower number of nodes. To define a dimensionality reduction procedure which is topologically informed, we need to estimate the topological properties of the spatial graphs. We leverage tools from Topological Data Analysis (TDA) [chazal2021introduction]. Taking its root in the field of algebraic topology, TDA is a recent and promising domain that provides well-founded tools to analyze complex data - such as graphs - for which topological properties represent key insights into understanding the data. Here, by topology we refer to the properties left unchanged by continuous deformations such as rotation, translation, stretching, and bending (e.g., number of connected components or holes). A common tool in TDA is the persistent diagram, which is classically used to study the topology of point clouds in $\mathbb{R}^{p}$ - see seminal works by [edelsbrunner2002topological, carlsson2009topology, aktas2019persistence]. As for point clouds, some attempt have been done to adapt classic filtration to graph data [archambault2007topolayout, li2012effective, ferrara2011topological]. For example, a graph filtration can be defined based on the node degree [hofer2017deep], the Jaccard index and Ricci curvature [zhao2019learning], or the Heat Kernel [carriere2020perslay, lasalle2024heat]. As our graphs have a spatial component, we define a new filtration which is based on both the spatial and the structural information. We name it the triangle-aware graph filtration as it captures cycles information generated by all cycles, even the one induced by triangles in the graph. Given the topological characterization, we can now propose a coarsening procedure for spatial graphs. The new graph preserve the most important topological features of the original graph - captured by our filtration and persistent diagram - with a lower number of nodes and edges. As for any dimensionality reduction technique, we would like to calibrate the reduction level, as too much reduction would yield to an oversimplified graph with missing important features, and too little reduction would not provide a significant and useful coarsening. We define a score to guide the choice of an appropriate reduction level. This score balances the reduced graph complexity with the amount of topological “distortion” between the original and reduced graph evaluated by comparing persistence diagrams. Minimizing this score allows to calibrate the reduction level.
@@
The paper is organized as follow. In Section 2, we introduce the spatial graph object and the coarsening procedure. In Section 3, we recall how to compute classic persistence diagrams and we introduce the persistence diagram adaptation for spatial graphs (Subsection 3.2). In Section 4, we define the topologically informed spatial graph coarsening. As the new methodology is tailored for spatial graphs, we study the invariance and equivariance properties of the method with respect to rotation, translation and scaling of the node positions; see Section 5. In Section 6, we showcase the proposed method on simulated data and on a real world dataset of fungi networks [fricker2025fungi]. The Python code of our method is publicly available in GitHub [github_repo].
@@ 2 Spatial Graph Coarsening
A spatial graph $G$ is given by a set of nodes $V=\{v_{1},\dots,v_{n}\}$, a set of edges $E\in V\times V$ and spatial positions of the nodes $(X_{u})_{u\in V}$, where $x_{u}\in\mathbb{R}^{p}$ for all nodes $u$, with $p$ being the space dimension. The spatial graph is denoted by the triplet $G=(V,E,X)$. In this work, the nodes are embedded in the space $\mathbb{R}^{p}$ equipped with standard Euclidean norm $\|\cdot\|$. In classic applications, the dimension $p$ is usually 2 or 3. The length function $\ell:E\rightarrow\mathbb{R}_{+}$ assigns to each edge $(u,v)\in E$ its edge length $\ell_{u,v}=d(x_{u},x_{v})$, computed via Euclidean distance.
The objective of this paper is to define a coarsening procedure for spatial graphs, retaining the main topological properties of the original graph. In a general graph setting, given a set of graphs $\mathcal{G}$, a coarsening method can be defined via a coarsening function $f:\mathcal{G}\rightarrow\mathcal{G}$ which defines a new graph on a partition of the nodes of the original graph: $f(G)=(f^{V}(G),f^{E}(G))$ where $f^{V}(G)\in\mathcal{P}(V)$, with $\mathcal{P}(V)$ denoted the set of partitions of $V$. As our graphs are spatial, we define a spatial graph coarsening methods based on a spatial parameter $\theta\geq 0$. Such parametrization of the coarsening procedure allows to define a coarsened graph which depends on both the structural and the spatial structure of the graph. Essentially, we obtain the coarsened spatial graph by collapsing the edges with length smaller than $\theta$. Before giving the formal definition of our spatial graph coarsening (Definition 1), we introduce the notation $sub(G,\theta)=(V,E_{\theta})$, such that $E_{\theta}=\{(u,v)\in E,\ s.t.\ \ell_{u,v}\leq\theta\}$. $sub(G,\theta)$ denotes the abstract subgraph of $G$, induced by the edges of length smaller than $\theta$.
@@
Consider a spatial graph $G=(V,E,X)$ and a parameter $\theta\geq 0$. We define the coarsened graph at scale $\theta$ by $f_{\theta}(G)=(f_{\theta}^{V}(G),f_{\theta}^{E}(G),f_{\theta}^{X}(G))$ where the three components $f_{\theta}^{V}(G)$, $f_{\theta}^{E}(G)$ and $f_{\theta}^{X}(G))$, acting respectively on the nodes, edges and positions, are defined as follows:
@@
    • the average positioning assigns to an hypernode $V_{i}$ the average position of the merged nodes:
@@
    • the degree positioning assigns to an hypernode $V_{i}$ the position of the node with the highest degree among the merged nodes:
@@
      where the degree $deg(v)\in\mathbb{R}_{+}$ is the number of edges incident to $v$.
@@
The parameter $\theta$ takes values in $\mathbb{R}_{+}$. However, in practice, the interesting set of values for the parameter is $\Theta=[\min_{(u,v)\in E}\ell_{u,v},\max_{(u,v)\in E}\ell_{u,v}]$. The coarsened graph for $\theta>\max_{(u,v)\in E}\ell_{u,v}$ is the trivial graph where all the nodes are collapsed into one single node. To simplify notations, we continue to consider $\theta$ in $\mathbb{R}_{+}$
Once we have introduced a spatial coarsening procedure, the question is now how to select the parameter $\theta$. The objective of the paper is to define a coarsened graphs with lower number of nodes while preserving the key topological features of the original graph. An illustrative example is shown in Figure 1. We sample $20$ nodes at random with position on an annulus shape. The corresponding fully connected graph has $(n^{2}-n)/2$ edges, but we select only a percentage of the edges, keeping only those with smallest length. The original graph is reduced into three graphs with lower number of nodes, obtained with different values for the parameter $\theta$. From the picture it is clear to see how the coarsening procedure is first merging the nodes while preserving its key topological feature which is largest the cycle. As the parameter $\theta$ increases, more nodes are merged together, resulting in an oversimplified graph. To identify a parameter $\theta$ that performs a good balance between reducing the graph and preserving the main topological features, we define a procedure that relies on persistence diagrams to capture the topological information.
@@ 3 Persistence Diagrams for Spatial Graphs
classic TDA pipelines consist in capturing certain properties for shapes underlying the data such as number of holes or number of connected components. One of the main tools is a topological descriptor called persistence diagram. In this section, we briefly introduce the general TDA pipeline for analyzing point clouds, before describing how they can be adapted to graph, and especially spatial graphs. For a more complete presentation of TDA, we refer the reader to [edelsbrunner2010computational],[oudot2015persistence], [chazal2021introduction], and [iniesta2022topological].
3.1 classic persistence diagrams
@@
Data as point clouds have no intrinsic interesting topology (only isolated contractible connected components). However, there often exist underlying continuous structures with richer topology, e.g., points sampled on some manifolds. To capture the topological invariants of the underlying shape, we construct continuous spaces on top of the point clouds called simplicial complexes. A simplicial complex is a space made by gluing together simple structures of different dimensions: points, segments, triangles, tetrahedrons etc. The simple building blocs are called simplices. They are represented in an abstract way as a set of their vertices, e.g., an segment between point $u$ and point $v$ is encoded as $\{u,v\}$, a triangle between $u$, $v$ and $w$ is given by $\{u,v,w\}$. Simplicial complexes are a generalization of classic graphs, which are one-dimensional objects (composed of points and edges only), where one can include higher-dimensional objects like triangles (dimension 2), tetrahedron (dimension 3) and equivalents in higher dimensions. Formally, they are a set of simplices, i.e., a set of sets of vertices. Given a point clouds $\mathcal{X}=\{x_{1},\dots x_{n}\}$ in $\mathbb{R}^{p}$ equipped with the Euclidean distance $d$, a classic simplicial complex built on top of $\mathcal{X}$ is the so-called Vietoris-Rips (VR) complex at scale $r>0$ denoted by $VR_{r}(\mathcal{X},d)$ [zomorodian2010fast]. To determine which simplices are included in $VR_{r}(\mathcal{X})$ consider all balls of radius $r$ centered on the points of $\mathcal{X}$. If two balls intersect, the corresponding edge between the two centers is an element of $VR_{r}(\mathcal{X},d)$. More generally, for higher dimension, any $k$-dimensional simplex (subset of $k+1$ points of $\mathcal{X}$) is included if all pairs of balls centered in points of the simplex intersect. Formally, the $k$-simplex $\{x_{i_{1}},\dots,x_{i_{k}}\}$ is in $VR_{r}(\mathcal{X},d)$, if for all $j$ and $j^{\prime}$ such that $1\leq j<j^{\prime}\leq k$, it satisfies $d(x_{i_{j}},x_{i_{j^{\prime}}})\leq r$.
@@
From this continuous structures one can now compute topological invariants of the data [chazal2021introduction]. To extract meaningful topological information from the data, instead of looking at the topological invariants of one simplicial complex at a given scale $r>0$, we can look at their evolution in the family of simplicial complexes $(VR_{r}(\mathcal{X},d))_{r>0}$ as $r$ increases. Such a family is called a filtration, i.e., a family of nested simplicial complexes. As $r$ increases, some topological features (connected components, holes, cavities, etc.) appear and disappear. For each topological feature, we record the scale $r_{b}$ at which it appears (it is born) and the scale $r_{d}\geq r_{b}$ at which it disappears (it dies). We encode this information into a point $(r_{b},r_{d})$ in $\mathbb{R}^{2}$. The collection of such points for all topological features is called the persistence diagram (PD). Points that are closer to the diagonal $y=x$ in the PD corresponds to topological features that lived during a short range of scales, they did not persist, hence they are usually associated to topological noise. Conversely, points that lie far from the diagonal correspond to topological features that persisted in the filtration and therefore correspond to important topological characteristics of the data. Note that monitoring the topological invariants in the filtration avoids having to choose a specific scale at which relevant topological features would be present. How much a topological feature persists in the filtration encode its relevance as a topological information of the data.